OpenStudy (anonymous):

F(x) = 3x-5; g(f(x)) = x; what is g(x)?

OpenStudy (anonymous):

g( f(x) ) means you are going to interpret g(x), where you replace x with the equation for f(x)

OpenStudy (anonymous):

Yes I am aware of that...

OpenStudy (anonymous):

g(3x-5) = X then I am stuck

OpenStudy (anonymous):

f(x) 3x-5. g( f(x) ) = x In This case, g(x) is the inverse function of f(x). Function with variable x into its inverse function will give you back x

OpenStudy (anonymous):

g(x) = (x+5)/3

OpenStudy (anonymous):

Take f(x) and do the inverse of all the operations in the correct order.

OpenStudy (anonymous):

I wanted to understand the mechanics of finding g(x) = x+5/3 A priori, I do not see that the two functions are inverse even though I know it.

OpenStudy (anonymous):

f(x)=3x-5. You divide by 3 first to get (x-5)/3. The opposite of subtracting is adding. (x+5)/3 = g(x)

OpenStudy (anonymous):

Please Excuse My Dear Aunt Sally Brackets, Exponents, Multiply/Divide, Add/Subtract in that order

OpenStudy (anonymous):

sorry, I did not mean that by mechanics Let me elaborate: g(3x-5) = X by replacing the f(x) value but after that I am lost in the algebraic mechanics Do you automatically see that fx and gx are inverse functions? and move from there?